# Balancing LoRA Performance and Efficiency with Simple Shard Sharing

### v10

@@ 5 Experiments
In this section, we evaluate the performance of MiSS on various benchmark datasets. Initially, we assess Natural Language Understanding (NLU) capabilities using a subset of the GLUE dataset with the robert-base model. Subsequently, we evaluated the Natural Language Generation (NLG [^2] Evaluations were performed with the OpenCompass repository, and the Math dataset was evaluated using a 5-shot prompt configuration.) capabilities by fine-tuning the LLM. To ensure fair comparisons, we select methods officially included in the PEFT repository to avoid discrepancies caused by inconsistent tuning implementations.
@@
Table 3: We conducted fine-tuning of large language models using MiSS and multiple LoRA variants, evaluating their performance on GSM8K, Math, HumanEval, and MBPP benchmarks. All reported results are averaged over three independent runs to ensure robustness. The first-place entry should be highlighted in bold, and the second-place entry should be underlined.
Model Strategy Trainable GSM8K [5] Math [26] HumanEval [4] Mbpp [1]
Llama2-7B [18] LoRA 89.9M 40.75 5.22 17.74 35.15
@@ 5.2 Experiment on Natural Language Generation
To verify the generalizability of MiSS, we conducted more comprehensive experiments on LLM. we conducted 3 more task finetuning experiments on LLM: math and code. (1) Math: We trained our model on a 395k subset of MetaMathQA [26], a dataset bootstrapped from other math instruction tuning datasets like GSM8K [5] and MATH  [26], with higher complexity and diversity. (2) Code: We train our model on a 100k subset of CodeFeedback [27], a high-quality code instruction dataset, removing explanations after code blocks. The model is tested on HumanEval [4] and Mbpp [1]. The hyperparameter settings for this experiment were kept equal, while the train steps were adjusted according to the specific fine-tuning datasets used. It is worth noting that the attention-based architectures employed by models such as LLaMA, Qwen, and Mistral do not use fully symmetric weight structures, which makes it impossible to achieve exact alignment of trainable parameters when comparing MiSS with LoRA. To address this, we set the rank $r$ of LoRA to 36 and the rank $r$ of MiSS to 64, ensuring that MiSS uses fewer parameters than LoRA to demonstrate its superiority. Each experiment is conducted with 2 different random seeds, and the average performance is reported.
32641282560612182430rankgradient normftLoRAPiSSAMiSS
Figure 4: Left. The loss curve for LLaMA2-7B fine-tuned on MetaMathQA shows MiSS’s superior fitting ability across architectures and parameter settings, with a notably rapid loss reduction in the first 100 steps, underscoring its effectiveness. Right. The initial gradient norm of model training.
@@ Results.
As shown in Table 6, MiSS consistently achieves state-of-the-art performance across multiple models and evaluation metrics. Specifically, on LLaMA2-7B, MiSS surpasses PiSSA on most metrics while using fewer trainable parameters than both LoRA and DoRA, showcasing its parameter efficiency and superior performance. On RWKV6-7B, where all methods share the same number of trainable parameters, MiSS still achieves the best overall performance, indicating its strong adaptation capability to non-transformer architectures. Across all evaluated models, including Mistral-7B, LLaMA2-13B, and Qwen3-4B, MiSS maintains its leading position on key benchmarks, often achieving the highest scores with significantly fewer parameters—further highlighting its effectiveness and scalability.
@@ 5.3 Effect of Rank $r$
This subsection explores the upper limits of the MiSS structure by varying the rank $r$ in the MiSS matrix. Comparative experiments were conducted by fine-tuning LLaMA2-7B  [^3] We use LLaMA2-7B instead of LLaMA3-8B because we found that the LLaMA3 series is over-optimized on math-related tasks. on the MetaMathQA dataset and validating on GSM8K and Math benchmarks. The test results, as shown in Table  4, demonstrate that the fine-tuning performance improves as the value of b increases. Notably, when $r=16$, the MiSS structure, with only one-quarter of the trainable parameters compared to PiSSA, surpasses PiSSA’s performance on the GSM8k benchmark. However, its performance on the Math benchmark is only 3.73. The GSM8K score surpasses that of PiSSA, but the Math score is significantly lower, indicating The size of $r$ impacts the model’s ability to understand unseen data. Based on this observation, we hypothesize that when the rank is too small, it significantly limits the model’s generalization ability.
@@
Figure 5: Time and space complexity. e denotes efficient form.
@@
5.4 Gradient Norm Analysis
To further investigate the adaptive capabilities of MiSS, we analyzed its initial gradient norms in comparison to Full Fine-tuning (FT), LoRA, and PiSSA. The magnitude of initial gradient norms is often correlated with faster convergence and the ability of a model to effectively adapt during fine-tuning. Our comparative analysis, illustrated in Figure 4 (Right), reveals distinct behaviors among the evaluated methods. Standard LoRA consistently exhibits the lowest initial gradient norms across the tested ranks. While PiSSA’s gradient norms increase with higher ranks, they generally remain significantly lower than those achieved by full fine-tuning. Notably, MiSS demonstrates substantially larger initial gradient norms compared to both LoRA and PiSSA across various ranks. More importantly, the gradient norm profile of MiSS closely approximates that of Full Fine-tuning. This proximity to FT’s gradient characteristics suggests that MiSS is capable of inducing more significant initial updates, potentially leading to a training dynamic more akin to full fine-tuning.
5.5 Resource and efficiency
Table  5 compares the training resources and token throughput required for fine-tuning RWKV6 using LoRA and MiSSon a single 4090 GPU. The specific fine-tuning settings are as follows: batch size = 1, context length (ctx_len) = 512. The results show that MiSS has the highest computational efficiency, being nearly 10% faster than LoRA while also being more memory-efficient. At the end of the table, we provide the actual resource costs for fine-tuning RWKV6 on the MetaMathQA dataset using 4 NVIDIA 4090 GPUs, with checkpoint techniques applied.
Table 5: Resource and efficiency
Devices Strategy Trainable Memory throughput
GPU $\times$ 1 LoRA 55.1M 12074 MB 3.62 kt/s
MiSS 55.1M 11052 MB 3.99 kt/s
GPU $\times$ 4 LoRA 55.1M 4$\times$15328 MB 15.6 kt/s
MiSS 55.1M 4$\times$15304 MB 16.0 kt/s
@@ 6 Conclusion
In this work, we proposed the Foissl framework, which divides pre-trained weights into multiple shards and updates them using a shared trainable matrix. This approach significantly reduces resource overhead and opens up new directions for PEFT techniques. Furthermore, we present a hardware-optimized version of MiSS, achieving remarkable gains in computational efficiency. Extensive experiments demonstrated the superiority of MiSS, which outperforms LoRA and its variants in evaluation metrics, computational efficiency, and resource usage.
Collectively, our framework redefines the design principles of efficient adaptation: MiSS provides theoretical grounding for dimension-wise decomposition, MiSS delivers practical efficiency parity with LoRA. This progression demonstrates that parameter efficiency need not come at the cost of expressivity, paving the way for adaptive fine-tuning in resource-constrained environments.
@@ 7 Limitations and Future work
Due to hardware constraints, we were unable to conduct and report results from the full-scale training that this method ideally requires. Nonetheless, through extensive evaluations and thorough comparisons with LoRA and its variants, we have validated the effectiveness of MiSS. The strong performance achieved on large language models gives us substantial confidence in the potential of this approach. We believe that MiSS can be readily applied to a wide range of multimodal tasks.
@@ Appendix B Settings of Experiments in No Free Lunch
Input $x$ Range $[-1,1]$
Training Samples ($N\_TRAIN$) 50
Validation Samples ($N\_VALID$) 100

### v11

@@ 5 Experiments
In this section, we evaluate the performance of MiSS on various benchmark datasets. Initially, we assess Natural Language Understanding (NLU) capabilities using a subset of the GLUE dataset . Subsequently, we evaluated the Natural Language Generation (NLG [^2] Evaluations were performed with the OpenCompass repository, and the Math dataset was evaluated using a 5-shot prompt configuration.) capabilities by fine-tuning the LLM.
@@
Table 3: We conduct a systematic comparison of LoRA, DoRA, PiSSA, and MiSS across several mainstream large language models (Llama2, RWKV, Mistral, and Qwen3). All reported results are averaged over three independent runs to ensure robustness. The first-place entry should be highlighted in bold, and the second-place entry should be underlined.
Model Strategy Trainable GSM8K Math HumanEval Mbpp
Llama2-7B (Touvron et al., 2023) LoRA 89.9M 40.75 5.22 17.74 35.15
@@ 5.2 Experiment on Natural Language Generation
To verify the generalizability of MiSS, we conducted more comprehensive experiments on LLM. we conducted 3 more task finetuning experiments on LLM: math and code. (1) Math: We trained our model on a 395k subset of MetaMathQA (Yu et al., 2023), a dataset bootstrapped from other math instruction tuning datasets like GSM8K (Cobbe et al., 2021) and MATH  (Yu et al., 2023), with higher complexity and diversity. (2) Code: We train our model on a 100k subset of CodeFeedback (Zheng et al., 2024), a high-quality code instruction dataset, removing explanations after code blocks. The model is tested on HumanEval (Chen et al., 2021) and Mbpp (Austin et al., 2021). The hyperparameter settings for this experiment were kept equal, while the train steps were adjusted according to the specific fine-tuning datasets used. It is worth noting that the attention-based architectures employed by models such as LLaMA, Qwen, and Mistral do not use fully symmetric weight structures, which makes it impossible to achieve exact alignment of trainable parameters when comparing MiSS with LoRA. To address this, we set the rank $r$ of LoRA to 36 and the rank $r$ of MiSS to 64, ensuring that MiSS uses fewer parameters than LoRA to demonstrate its superiority. Each experiment is conducted with 2 different random seeds, and the average performance is reported.
@@ Results.
(a) Loss-Token
(b) Loss-Time
Figure 4: Loss curves of LLaMA2-7B fine-tuned on MetaMathQA using LoRA and MiSS(̇a) Loss vs. tokens. (b) Loss vs. training time.
The results (See Table  3) show that MiSS consistently achieves the best or near-best performance across different model scales and diverse tasks. For example, on Llama2-7B, MiSS significantly surpasses LoRA on GSM8K (48.16 vs. 40.75) and also improves HumanEval and Mbpp with scores of 8.58 and 36.81, respectively. As the model scales up to Llama2-13B, MiSS maintains its advantage, particularly on code generation benchmarks such as HumanEval and Mbpp, outperforming both PiSSA and DoRA. More importantly, on the latest large-scale model Qwen3-4B, which is trained on richer data, MiSS achieves the highest Math score of 34.82, far exceeding other methods, while remaining competitive on GSM8K and Mbpp. These findings demonstrate that MiSS is not only effective on medium-sized models but also scales robustly to larger architectures and data-rich models.
@@ 5.3 Effect of Rank $r$
We evaluate MiSS with varying matrix ranks to study the trade-off between tuning capacity and parameter cost. The Table 4 reports results for ranks $r\in\{16,32,64,128\}$ (corresponding to $\{21.7\text{M},43.5\text{M},87.0\text{M},174.0\text{M}\}$ trainable parameters). Performance on GSM8K and the Math benchmark improves monotonically as the rank increases: GSM8K rises from 45.90 at $r=16$ to 53.49 at $r=128$, while Math increases from 3.77 to 10.08. In practice, $r=64$ offers a favorable trade-off (48.16 GSM8K, 8.58 Math) between performance gains and parameter overhead.
@@
Table 5: Time and space complexity. e denotes efficient form.
@@
5.4 Pareto plot
The emergence of PEFT techniques is motivated by dual objectives: mitigating GPU memory constraints and exploring more efficient model architectures. Nevertheless, numerous contemporary studies disproportionately focus on ultimate performance benchmarks, overlooking critical practical considerations like computational efficiency and training duration—an emphasis that clearly diverges from the original rationale for PEFT. In this section, we undertake a multi-dimensional investigation into the relationships among computational overhead, efficiency, and performance for diverse models. Leveraging the official Hugging Face PEFT (Mangrulkar et al., 2022) benchmarking framework, our evaluations are conducted under fair and reproducible conditions.
The Pareto frontiers in our evaluation provide definitive evidence of MiSS’s effectiveness. In every experimental setting, MiSS is uniquely positioned in the top-left corner—the optimal region—signifying that it delivers the best performance with minimal efficiency cost. This consistent advantage underscores MiSS’s unique contribution in balancing these competing objectives.
Figure 5: Pareto front of MiSS comparing with other PEFT methods. We select three more methods as the baseline on the balancing of memory and performance.
Table 6: Experimental results across PEFT methods on Llama-3.2-3B.
PEFT Type Total Time Train Time Test Accuracy Train Loss Accelerator Memory (Bytes)
Max Reserved 99th Reserved Avg
RSLORA 2069 1871 0.5299 0.5657 22,538,092,544 17,953,927,987 12,128,059,444
C3A 2125 1924 0.5102 0.5808 22,280,142,848 17,825,917,829 11,804,454,210
[…]
RANDLORA 2457 2213 0.5072 0.5785 22,798,139,392 18,436,063,232 12,743,670,025
SHIRA 2085 1867 0.5072 0.5789 21,743,271,936 17,637,383,864 12,240,924,809
OFT 2494 2214 0.5057 0.5947 22,294,822,912 17,939,310,837 12,057,354,384
LORA 1993 1796 0.4822 0.6069 22,273,851,392 17,710,763,212 11,868,689,976
DORA 2287 2023 0.4807 0.6068 24,553,455,616 19,189,150,515 12,490,471,636
LORAFA 2026 1821 0.4299 0.6510 20,187,185,152 16,257,394,933 11,106,307,276
LOHA 2591 2341 0.4185 0.6570 23,886,561,280 19,247,870,771 13,446,820,344
IA3 1922 1746 0.4124 0.6569 23,135,780,864 18,398,356,439 12,023,331,867
ADALORA 2209 1986 0.3904 0.6863 22,793,945,088 18,203,426,160 12,361,399,900
LOKR 2352 2152 0.3753 0.6877 23,565,697,024 18,987,698,094 13,173,683,073
P_TUNING 1918 1707 0.3707 0.6740 20,937,965,568 17,215,688,540 11,867,101,593
VBLORA 2210 1962 0.3700 0.7143 22,181,576,704 17,635,223,797 11,735,344,663
VERA 2025 1820 0.3685 0.6927 21,596,471,296 17,291,123,097 11,489,715,316
BOFT 11,114 8292 0.3647 0.7268 24,427,626,496 20,103,445,872 14,814,855,089
IA3 2005 1783 0.3450 0.7657 23,137,878,016 18,398,566,154 12,023,227,429
TRAINABLE_TOKENS 1814 1572 0.2881 0.7862 20,956,839,936 16,957,675,929 12,730,137,942
PROMPT_TUNING 2715 2394 0.2525 0.7790 24,408,752,128 20,650,676,715 15,297,364,466
ADAPTION_PROMPT 2261 1989 0.2206 0.8317 22,410,166,272 17,907,664,814 11,893,757,234
PREFIX_TUNING 1959 1662 0.1471 0.7887 20,912,799,744 16,945,051,074 11,766,684,083
FOURIERFT 2824 2422 0.1198 0.9979 23,681,040,384 19,054,869,872 13,111,221,498
PROMPT_TUNING 2700 2380 0.0500 1.0655 24,379,392,000 20,669,781,770 15,297,773,830
FOURIERFT 2824 2424 0.0008 1.2480 23,653,777,408 19,017,267,937 13,104,129,350
LN_TUNING 1870 1657 0.0000 1.2370 21,177,040,896 16,903,066,091 11,385,589,622
@@ 6 Conclusion
This work tackles the critical inefficiency of simultaneous matrix updates in Low-Rank Adaptation (LoRA), which leads to slow convergence and suboptimal resource use. We propose MiSS as a compelling solution—a new PEFT framework that updates decomposed weight shards using a single, shared matrix. This approach drastically reduces optimization complexity and resource demands. Comprehensive experiments validate that MiSS consistently outperforms existing methods in accuracy, memory footprint, and computational speed, offering a fundamentally more efficient pathway for adapting large models.
@@ 7 Limitations and Future Work
Through extensive evaluations and thorough comparisons with LoRA and its variants, we have validated the effectiveness of MiSS. The strong performance achieved on large language models gives us substantial confidence in the potential of this approach. We believe that MiSS can be readily applied to a wide range of multimodal tasks.
@@ Appendix B Settings of Experiments in No Free Lunch
Input $x$ Range [1,1][-1,1]
Training Samples (N_TRAINN\_TRAIN) 50
Validation Samples (N_VALIDN\_VALID) 100